# Plasmon-plasmon interaction in nanoparticle assemblies: role of the dipole-quadrupole coupling

### v1

@@ IV Conclusion and perspectives
Having developed a quasistatic Hamiltonian (2) from the plasmonic overlap of classical and the quantum models for plasmonic arrays, we have shown that dipolar and quadrupolar modes represent the major contribution of the low-energy physics and can be decoupled from higher-order modes. Consequently, we proposed a minimal model that treats dipoles and quadrupoles on the same footing and allows an accurate description of the two first plasmonic excitations across the whole Brillouin zone of a chain of nanoparticles. Further inspection of the nature of the modes underlined the importance of dark quadrupoles in dense particle arrays, solely due to plasmon-plasmon interaction. We pointed out that even though quadrupoles do not couple to dipoles at high symmetry points ($k=0$ for the chain), they do at others, rendering them essential to investigate plasmon-polaritons in nanoparticle assemblies. The coupling of the present Hamiltonian to a photonic bath $H_{\gamma}$ that would allow taking account of retardation effects and computing physical observables is left for further investigation.

### v2

@@ IV Conclusion and perspectives
Having developed a quasistatic Hamiltonian (2) from the plasmonic overlap of classical and the quantum models for plasmonic arrays, we have shown that dipolar and quadrupolar modes represent the major contribution of the low-energy physics and can be decoupled from higher-order modes. Consequently, we proposed a minimal model that treats dipoles and quadrupoles on the same footing and allows an accurate description of the two first plasmonic excitations across the whole Brillouin zone of a chain of nanoparticles. Further inspection of the nature of the modes underlined the importance of quadrupoles in dense particle arrays, solely due to plasmon-plasmon interactions. For this reason, in future work considering also the coupling with light, the quadrupole will have to be accounted for regardless of how little they couple to light. We pointed out that even though quadrupoles do not couple to dipoles at high symmetry points ($k=0$ for the chain), they do at others, rendering them essential to investigate plasmon-polaritons in nanoparticle assemblies. The coupling of the present Hamiltonian to a photonic bath $H_{\gamma}$ that would allow taking account of retardation effects and computing physical observables is left for further investigation. Finally, the second quantized formalism used here makes our Hamiltonian relevant for coupling to other systems, such as molecules or quantum dots, that couple to the highly localized fields of quadrupoles.